Suppose that business travelers and vacationers

have the following demand for airline tickets

from New York to Boston:

Quantity Demanded Quantity Demanded

Price (business travelers) (vacationers)

$150 2,100 tickets 1,000 tickets

200 2,000 800

250 1,900 600

300 1,800 400

a. As the price of tickets rises from $200 to

$250, what is the price elasticity of demand

for (i) business travelers and (ii) vacationers?

(Use the midpoint method in your

calculations.)

b. Why might vacationers have a different

elasticity from business travelers?


1
Expert's answer
2021-10-04T16:11:43-0400

Elasticity of demand depicts how much a consumer responds to change in price level.

(a)

"Ed=\frac{\frac{Q_2-Q_1}{\frac{Q_2+Q_1}{2}}}{\frac{P_2-P_1}{\frac{P_2+P_1}{2}}}"


(i)

For business travelers


"Ped,Ed=\frac{ \frac{1900-2000}{\frac{1900+2000}{2}}}{\frac{250-200}{\frac{250+200}{2}}}"


"Ed=-0.23"

"|Ed|=-0.23<1; inelastic"



(ii)

"Ped,Ed=\frac{ \frac{600-800}{\frac{600+800}{2}}}{\frac{250-200}{\frac{250+200}{2}}}"



"Ed=-1.29"

"|Ed|=-1.29<1;inelastic"


(b)

Vacationers are casual with their plan whereas business travelers are traveling with business purpose thus their elasticity of demand for tickets is less affected with change in price.